# Entanglement and apparent thermality in simulated black holes

### v1

@@ IV Entanglement entropy of chiral spin-chain simulator
In the previous section, we derived analytic expressions for the fermion zero-mode entropy of a black hole space-time’s sub-region, under the assumption that the interior region is in thermal equilibrium with the exterior region. We now numerically compute the entanglement entropy of the chiral spin-chain for arbitrary bipartitions. We do this for two distinct cases: first, when the partition is located at the event horizon; and second, when the partition is located at a lattice site in the region corresponding to the black hole’s interior or exterior. We expect that the fermion zero-mode entropy agrees with the entanglement entropy if the ground state of the partitioned system is thermal, as, in this case, the von Neumann entropy equates to the thermodynamic entropy of the subsystem’s microstates (fermion zero-modes) [13, 36]. Thus, by comparing the agreement between our analytic predictions and numerical results, we can identify the regimes where this equality and, by extension, the assumption of thermality is valid and where it breaks down. We will see that as the system is free, the thermality condition is generally not valid, as, due to the absence of interactions, the system cannot truly thermalise. Surprisingly, this equality only holds when partitioning at the event horizon, due to the unique physics emerging in that region [5, 6].
@@
Consider a quantum system that is bipartitioned into two subsystems, $\mathcal{A}$ and $\mathcal{B}$, as illustrated in Fig. 1a, with a ground state $|\psi_{\mathcal{AB}}\rangle\in\mathcal{H_{A}}\otimes\mathcal{H_{B}}$ can described by the density matrix $\rho_{\mathcal{AB}}=|\psi_{\mathcal{AB}}\rangle\langle\psi_{\mathcal{AB}}|$, where $\mathcal{H_{A}}$ and $\mathcal{H_{B}}$ denote the Hilbert spaces of $\mathcal{A}$ and $\mathcal{B}$, respectively. The bipartite entanglement entropy $S_{\mathcal{A}}$ of subsystem $\mathcal{A}$, which quantifies the correlations between $\mathcal{A}$ and $\mathcal{B}$, is defined as $S_{\mathcal{A}}=-\text{Tr}\left(\rho_{\mathcal{A}}\ln(\rho_{\mathcal{A}})\right)$, where $\rho_{\mathcal{A}}=\text{Tr}_{\mathcal{B}}(\rho_{\mathcal{AB}})=\sum_{i}\langle i_{\mathcal{B}}|\rho_{\mathcal{AB}}|i_{\mathcal{B}}\rangle$ denotes the reduced density matrix of $\mathcal{A}$ and $\{|i_{\mathcal{B}}\rangle\}\in\mathcal{H_{B}}$ the orthonormal basis of $\mathcal{H_{B}}$.
@@
Figure 1: $(a)$ Schematic of spin-chain bipartition at lattice site $n_{s}$ located a distance $\Delta n$ from $n_{h}$. $(b)$ Red data points show entanglement entropy $S_{\mathcal{A}}(n_{s})$ of $\mathcal{A}$ as a function of the partition site $n_{s}$ for a system with $n_{h}=1000$. Blue and orange curves show the fermion zero-mode entropies of Eqs.(19) and (22) for the black hole space-time’s interior and exterior regions, with $N_{F}=2$ and $\tilde{N}_{F}=1$, respectively, and the non-universal constants chosen such that the each curve reduces to $S_{\mathcal{A}}(n_{h})$ in the limit $n_{s}\rightarrow n_{h}\mp 2$, respectively. (c), (d), and (e) Red data points show the entanglement entropy $S_{\mathcal{A}}(n_{s})$ of $\mathcal{A}$ as a function of the partition’s location, where the system is partitioned at $n_{s}=n_{h}-\Delta n$ (interior), the event horizon $n_{h}$, and $n_{s}=n_{h}+\Delta n$ (exterior), respectively, with $\Delta n=250$. Blue curves show data interpolated using Eqs.(19), (20), and (22) for the fermion zero-modes’ entropy for the interior, horizon, and exterior regions of the black hole space-time, respectively. To avoid partitioning the two-site unit cell of the continuum limit, only even $n_{h}$ were considered. For all figures, systems of $N=10000$ lattice sites with couplings $v(n)=\sqrt{n_{h}/n}$ and $u=1$ were taken.
@@
We take the Hamiltonian of Eq. (2) with couplings $v(n)=\sqrt{n_{h}/n}$ and $u=1$, which represent the Schwarzchild black hole profile, where we have introduced the discrete spatial coordinate $n=x/2a$ and $a=2a_{c}=1/N$ is the chiral spin-chain’s lattice spacing. Moreover, $n_{h}$ is the lattice site where the event horizon is located, defined as the site for which $v(n_{h})=1$. As the Hamiltonian of Eq. (2) is that of a free fermion system, the entanglement entropy can be efficiently computed via
@@
We now turn to the analytic expressions for the entropy given in Eqs. (19), (20), and (22), which correspond to partitions in the interior, at the horizon, and in the exterior regions of the black hole’s space-time, respectively, where we parametrise the partition’s location by the integer $n_{s}=s/2a$ that is an absolute distance $\Delta n\geq 2$ from $n_{h}$. In these formulas, the integer parameters $N_{F}$ and $\tilde{N}_{F}$ remain to be determined. To fix these parameters, we numerically evaluate the entanglement entropy for a fixed $\Delta n$ as we vary the horizon’s position $n_{h}$ and, hence, the partition $n_{s}$, as shown in Figs. 1c, d, and e. We find that, up to an overall non-universal constant, only the expression corresponding to the horizon partition yields $N_{F}\approx 2$, with this value converging as $N\rightarrow\infty$, whilst the interior partition suggests a larger value and the exterior a smaller one, both being in general non-integers. The convergence of $N_{F}\rightarrow 2$ in the thermodynamic limit, $N\rightarrow\infty$, for the horizon partition can be inferred from the entanglement entropy’s derivative, which, as shown in Fig. 2, plateaus around the value 2 for larger domains as the system size increases. In contrast to this, the deviation of $N_{F}$ from an integer value as the partition moves from the horizon is emphasised also in Fig. 2, where we plot its value from fitting the analytic formulas as $n_{h}$ varies for different $\Delta n$; we see that $N_{F}$ is not a constant, in contradiction to the assumptions used to determine the fermion zero-mode entropy.
@@
To illustrate the entropy for general partitions $n_{s}\neq n_{h}$, we adopt $N_{F}=2$ for the interior and $\tilde{N}_{F}=1$ for the exterior, reflecting the central charge values known for homogeneous couplings [25, 26]. Moreover, we fix the non-universal constants such that the analytic entropies match the numerical results when the partition is located at $n_{h}$, where the numerically determined entanglement entropy and the theoretically derived black hole fermion zero-mode entropy exhibit the same scaling behaviour, yielding the expected integer value $N_{F}=2$. Nevertheless, as illustrated in Fig. 1b, we observe that these analytic entropy formulas differ significantly from the numerically computed entanglement entropy in both the interior and exterior. Under the assumption of thermality, we expect the fermion zero-mode entropy to be equivalent to the entanglement entropy between the partitioned subsystems [13, 36]; thus, their increasing discrepancy as the partition deviates from the event horizon signals the failure of thermality in the system for general partitions. We also attribute the lack of consistent integer solutions for $N_{F}$ and $\tilde{N}_{F}$ to this breakdown of assumed global thermality that underpins the analytic derivations. In Sec. V, this is verified by studying the subsystem’s mode occupation, which reveals a Fermi-Dirac distribution only when the lattice is partitioned near the event horizon.
@@
Figure 2: Data points show the interpolated value for the number of fermion species $N_{F}$ as a function of the distance $\Delta n$ of the partition $n_{s}=n_{h}-\Delta n$ from the event horizon $n_{h}$, for a system of $N=10000$ lattice sites with couplings $v(n)=\sqrt{n_{h}/n}$ and $u=1$. Each interpolated value of $N_{F}$ is determined by fitting Eq. (19) to numerical data for the entanglement entropy for a fixed $\Delta n$ as $n_{h}$ is increased. The deviation from $N_{F}\approx 2$ as $\Delta n$ increases indicates the breakdown of Eq. (19) for the fermion zero-mode entropy of the black hole’s interior in describing the behaviour of the entanglement entropy as the partition deviates from the event horizon. Inset shows the derivative of the entanglement entropy $S_{\mathcal{A}}(n_{h})$ with respect to $\ln(2n_{h})$ following a horizon partition, $\Delta n=0$, for increasing system sizes, indicating the regime where $N_{F}\approx 2$. As system size increases, the domain for which the gradient plateaus around 2 increases, suggesting that $N_{F}\rightarrow 2$ in the thermodynamic limit, $N\rightarrow\infty$.
@@ V.1 The breakdown of thermality
To further investigate the breakdown of thermality for non-horizon partitions, we examine the mode occupation expectation value, $\langle 0_{M}|c_{k}^{\dagger}c_{k}|0_{M}\rangle$, of the Gullstrand-Painlevé Hamiltonian (Eq. (2) with renormalised couplings $v(x)=\sqrt{x_{h}/x}$ and $u=1$), computed with respect to the ground state $|0_{M}\rangle$ of the Minkowski Hamiltonian (Eq. (2) with renormalised couplings $v=0$ and $u=1$). Here, $c_{k}$ are the eigenmodes that diagonalise the mean-field Hamiltonian of Eq. (2) [25]. The mode occupation can be expressed in terms of the single particle Gullstrand-Painlevé and Minkowski Hamiltonians’ eigenstates $|E_{GP,k}\rangle$ and $|E_{M,q}\rangle$, respectively, as follows: [49]
@@
where $k$ and $q$ are indices that label the Gullstrand-Painlevé and Minkowski Hamiltonians’ eigenstates, respectively.
To perform the partial trace of subsystem $\mathcal{A}$, we define the Gullstrand-Painlevé Hamiltonian only in the region corresponding to subsystem $\mathcal{B}$ and take it to be zero elsewhere. This effectively traces out the degrees of freedom of $\mathcal{A}$. As illustrated in Fig. 3, when the system is partitioned at the event horizon ($\Delta n=0$), the mode occupation expectation value follows a Fermi-Dirac distribution of the form
$f(E_{k},T)=\frac{1}{e^{E_{k}/T}+1},$ (25)
@@
Figure 3: Red data points show the mode occupation expectation value $\langle 0_{M}|c_{k}^{\dagger}c_{k}|0_{M}\rangle$ of the Gullstrand-Painlevé Hamiltonian (Eq.(2) with renormalised couplings $v(x)=\sqrt{x_{h}/x}$ and $u=1$, defined only on $\mathcal{B}$) with respect to the ground state $|0_{M}\rangle$ of the Minkowski Hamiltonian (Eq. (2) with renormalised couplings $u=1$ and $v=0$, defined over the entire lattice) for a horizon partition ($\Delta n=0$) and interior partition ($\Delta n=50$), respectively. The $x$-axes show the Gullstrand-Painlevé Hamiltonian’s eigenvalues $E_{GP,k}$. Blue curve shows data interpolated using the Fermi-Dirac distribution of Eq. (25) with $T=0.3271$ and $T=1.9237$, respectively. System sizes of $N=2000$ lattice sites, with the horizon located at $n_{h}=700$, were considered.
Figure 4: (a) Absolute value of interpolation errors $|e(E_{GP,k})|$ between the numerical and interpolated data for the mode occupation number $\langle 0_{M}|c_{k}^{\dagger}c_{k}|0_{M}\rangle$, interpolated using a Fermi-Dirac distribution, for systems of increasing $\Delta n$ with $N=2000$ lattice sites. As $\Delta n$ increases, and the system is partitioned further from $n_{h}$, increasing absolute interpolation errors indicate deviations from a Fermi-Dirac distribution in the mode occupation expectation value. (b) Integrated absolute interpolation errors $I_{e}(E_{GP,k})$ for increasing $\Delta n$ and varied system sizes NN, illustrating that the deviations from a Fermi-Dirac distribution as $\Delta n$ is increased are not minimised at larger system sizes. This is in contrast to partitioning the system at $n_{h}$, where the integrated interpolation errors decrease for larger system sizes. For both figures, systems with $n_{h}=N/4$ were considered.
@@ V.2 Thermality and the Hawking temperature
We employ this method to probe the thermal properties of our chiral spin-chain simulator. However, the mean-field Hamiltonian of Eq. (2) does not feature an overall linear coupling and, therefore, does not approximate the entanglement Hamiltonian required to extract a thermal spectrum. To address this, the mean-field Hamiltonian must be modified such that the black hole geometry is encoded with only a linear nearest-neighbour coupling $u(x)$. In the continuum limit, such a modification corresponds to performing a coordinate transformation from Gullstrand–Painlevé to Schwarzschild coordinates, whilst preserving the space-time geometry; thus, by mapping the black hole geometry to a mean-field Hamiltonian that approximates the entanglement Hamiltonian, the mode occupations of the partitioned ground state yields a Fermi-Dirac distribution from which $T_{H}$ can be reliably extracted [49, 51]. This procedure fails, however, when the chirality term $v(x)$ is non-zero, as in this case the mean-field Hamiltonian no longer approximates the corresponding entanglement Hamiltonian.
@@
which matches the Schwarzschild metric in $(1+1)$ dimensions. In particular, for the choice $u(x)=1-x_{h}/x$, this line element describes the exterior region of a Schwarzschild black hole with an event horizon located at $x_{h}$. This way, taking $v(x)$ to be zero and making $u(x)$ position dependent corresponds to a coordinate transformation from Gullstrand–Painlevé coordinates $(t,x)$ to Schwarzschild coordinates $(\tilde{t},x)$. Whilst the Hamiltonian’s structure changes under this transformation, scalar quantities, such as the space-time curvature, remain invariant.
@@
According to the Fulling–Davies–Unruh effect [53, 54, 55], such an observer perceives the Minkowski vacuum as a thermal state with temperature $T_{H}=\kappa/2\pi$ [4] (see Appendices C.). It follows that the partitioned ground state of the mean-field Hamiltonian of Eq. (2), with the linear coupling $u(x)$ of Eq. (27) and $v=0$, will be thermal with an effective temperature equal to the Hawking temperature $T_{H}$, as it approximates both the entanglement Hamiltonian and the Hamiltonian of a Dirac spinor in Rindler space-time [49, 51].
As in the Gullstrand–Painlevé case, the thermality of the Hamiltonian corresponding to a Dirac spinor in Rindler space-time can be confirmed by analysing the mode occupation of the partitioned system’s ground state. Specifically, we compute the expectation value of the mode occupation operators with respect to the Minkowski vacuum state $|0_{M}\rangle$, corresponding to the ground state of the Minkowski Hamiltonian. As shown in Fig. 5, the resulting mode occupations for the Rindler Hamiltonian follow a Fermi–Dirac distribution with temperature $T_{H}=\kappa/2\pi$, thereby confirming the thermal nature of the reduced state and validating the extraction of the Hawking temperature via this approach. Once again, this thermality is only present when the lattice system is partitioned at the horizon.
Figure 5: Data points show the mode occupation expectation values $\langle 0_{M}|c_{k}^{\dagger}c_{k}|0_{M}\rangle$ of the Rindler Hamiltonian (Eq. (2) with renormalised couplings $u(x)=\alpha(x/x_{h}-1)$ and $v=0$, defined only on $\mathcal{B}$), computed with respect to the Minkowski ground state $|0_{M}\rangle$ (Eq. (1) with $u=1$ and $v=0$, defined over the entire lattice). The horizontal axis shows the eigenvalues $E_{R,k}$ of the Rindler Hamiltonian. The blue curves correspond to a Fermi–Dirac distribution with temperature $T_{H}=\kappa/2\pi$, where $\kappa=u^{\prime}(x_{h})/2=\alpha/2x_{h}$ is the surface gravity associated with the linear coupling profile. System sizes $N=2000$, $2400$, $2800$, $3000$, and $3600$ were considered, with the horizon located at $n_{h}=N/2$ and coupling slope $\alpha=2$. Inset depicts the absolute interpolation error $|e(E_{R,k})|$ between numerical and interpolated data of the mode occupation for $\Delta n=0,20,40,60,80$ (colour scheme as in Fig. 4a) with $N=2000$ and $\alpha=1/2$.

### v2

@@ IV Entanglement entropy of chiral spin-chain simulator
In the previous section, having assumed that the interior region is well described by a thermal state, we derived analytic expressions for the fermion zero-mode entropy for various regions of the black hole’s space-time. We now numerically compute the entanglement entropy of the chiral spin-chain for arbitrary bipartitions. We do this for two distinct cases: first, when the partition is located at the event horizon; and second, when the partition is located at a lattice site in the region corresponding to the black hole’s interior or exterior. We expect that the fermion zero-mode entropy agrees with the entanglement entropy if the ground state of the partitioned system is thermal, as, in this case, the von Neumann entropy equates to the thermodynamic entropy of the subsystem’s microstates (fermion zero-modes) [70, 42]. Thus, by comparing our analytic predictions and numerical results, we can identify the regimes where this equality and, by extension, the assumption of thermality is valid, and where it breaks down. We will see that as the system is free, the thermality condition is generally not valid, as, due to the absence of interactions, the system cannot truly thermalise. Surprisingly, this equality holds only when partitioning at the event horizon, due to the unique physics emerging in that region [27, 28].
@@
Consider a quantum system that is bipartitioned into two subsystems, $\mathcal{A}$ and $\mathcal{B}$, as illustrated in Fig. 3a, with a ground state $|\psi_{\mathcal{AB}}\rangle\in\mathcal{H_{A}}\otimes\mathcal{H_{B}}$ described by the density matrix $\rho_{\mathcal{AB}}=|\psi_{\mathcal{AB}}\rangle\langle\psi_{\mathcal{AB}}|$, where $\mathcal{H_{A}}$ and $\mathcal{H_{B}}$ denote the Hilbert spaces of $\mathcal{A}$ and $\mathcal{B}$, respectively. The bipartite entanglement entropy $S_{\mathcal{A}}$ of subsystem $\mathcal{A}$, which quantifies the correlations between $\mathcal{A}$ and $\mathcal{B}$, is defined as $S_{\mathcal{A}}=-\text{Tr}\left(\rho_{\mathcal{A}}\ln(\rho_{\mathcal{A}})\right)$, where $\rho_{\mathcal{A}}=\text{Tr}_{\mathcal{B}}(\rho_{\mathcal{AB}})=\sum_{i}\langle i_{\mathcal{B}}|\rho_{\mathcal{AB}}|i_{\mathcal{B}}\rangle$ denotes the reduced density matrix of $\mathcal{A}$ and $\{|i_{\mathcal{B}}\rangle\}\in\mathcal{H_{B}}$ the orthonormal basis of $\mathcal{H_{B}}$.
@@
Figure 3: $(a)$ Schematic of spin-chain bipartition at lattice site $n_{s}$ located a distance $\Delta n$ from $n_{h}$. $(b)$ Red data points show entanglement entropy $S_{\mathcal{A}}(n_{s})$ of $\mathcal{A}$ as a function of the partition site $n_{s}$ for a system with $n_{h}=1000$. Blue and orange curves show the fermion zero-mode entropies of Eqs. (18) and (21) for the black hole space-time’s interior and exterior regions, respectively, with $N_{F}=\tilde{N}_{F}=1$ and the non-universal constant of the exterior region chosen to match that obtained for the horizon partition. (c), (d), and (e) Red data points show the entanglement entropy $S_{\mathcal{A}}(n_{s})$ of $\mathcal{A}$ as a function of the partition’s location, where the system is partitioned at $n_{s}=n_{h}-\Delta n$ (interior), the event horizon $n_{h}$, and $n_{s}=n_{h}+\Delta n$ (exterior), respectively, with $\Delta n=250$. Blue curves show data interpolated using Eqs.(18), (19), and (21) for the fermion zero-modes’ entropy for the interior, horizon, and exterior regions of the black hole space-time, respectively. To avoid partitioning the two-site unit cell of the continuum limit, only even $n_{h}$ were considered. For all figures, systems of $N=10000$ lattice sites with couplings $v(n)=\sqrt{n_{h}/n}$ and $u=1$ were taken.
@@
We take the Hamiltonian of Eq. (2) with couplings $v(n)=\sqrt{n_{h}/n}$ and $u=1$, which represent the Schwarzchild black hole profile, where we have introduced the discrete spatial coordinate $n=x/2a$ and $a=a_{c}/2=1/N$ is the chiral spin-chain’s lattice spacing. Moreover, $n_{h}$ is the lattice site where the event horizon is located, defined as the site for which $v(n_{h})=1$. As the Hamiltonian of Eq. (2) is that of a free fermion system, the entanglement entropy can be efficiently computed via
@@
We now turn to the analytic expressions for the entropy given in Eqs. (18), (19), and (21), which correspond to partitions in the interior, at the horizon, and in the exterior regions of the black hole’s space-time, respectively, where we parametrise the partition’s location by the integer $n_{s}=s/2a$ that is an absolute distance $\Delta n\geq 2$ from $n_{h}$. In these formulas, the integer parameters $N_{F}$ and $\tilde{N}_{F}$ remain to be determined. To fix these parameters, we numerically evaluate the entanglement entropy for a fixed $\Delta n$ as we vary the horizon’s position $n_{h}$ and, hence, the partition $n_{s}$, as shown in Figs. 3c, d, and e. We find that, up to an overall additive constant, which is non-universal with respect to the chosen lattice parametrisation, only the expression corresponding to the horizon partition yields an integer value of $N_{F}\approx 1$, with this value converging as $N\rightarrow\infty$, whilst interior and exterior partitions suggest a larger value, both being in general non-integers. The convergence of $N_{F}\rightarrow 1$ in the thermodynamic limit, $N\rightarrow\infty$, for the horizon partition can be inferred from the entanglement entropy’s derivative, which, as shown in Fig. 4, plateaus around the value 1 for larger domains as the system size increases. In contrast to this, the deviation of $N_{F}$ from an integer value as the partition moves from the horizon is emphasised in Fig. 4, where we plot its value from fitting the analytic formulas as $n_{h}$ varies for different $\Delta n$; we see that $N_{F}$ is not a constant, in contradiction to the assumptions used to determine the fermion zero-mode entropy.
@@
To illustrate the entropy for general partitions $n_{s}\neq n_{h}$, we adopt $N_{F}=1$ for the interior and $\tilde{N}_{F}=1$ for the exterior, reflecting the value of the number of fermionic fields determined for a horizon partition. Moreover, we fix the non-universal constant of the exterior region’s analytic zero-mode entropy to match the value obtained when partitioning at the horizon. As illustrated in Fig. 3b, we observe that the analytic zero-mode entropy formulas differ significantly from the numerically computed entanglement entropy in both the interior and exterior. Under the assumption of thermality, we expect the fermion zero-mode entropy to be equivalent to the entanglement entropy between the partitioned subsystems [70, 42]; thus, the increasing discrepancy between them as the partition deviates from the horizon signals the failure of thermality for general partitions. We also attribute the lack of consistent integer solutions for $N_{F}$ and $\tilde{N}_{F}$ to this breakdown of the assumed thermality that underpins the analytic derivations. In Sec. V, this is verified by studying the subsystem’s mode occupation, which reveals a Fermi-Dirac distribution only when the lattice is partitioned near the event horizon.
@@
Figure 4: Data points show the interpolated value for the number of fermionic fields $N_{F}$ as a function of the distance $\Delta n$ of the partition $n_{s}=n_{h}-\Delta n$ from the event horizon $n_{h}$, for a system of $N=10000$ lattice sites with couplings $v(n)=\sqrt{n_{h}/n}$ and $u=1$. Each interpolated value of $N_{F}$ is determined by fitting Eq. (18) to numerical data for the entanglement entropy for a fixed $\Delta n$ as $n_{h}$ is increased. The deviation from $N_{F}\approx 1$ as $\Delta n$ increases indicates the breakdown of Eq. (18) for the fermion zero-mode entropy of the black hole’s interior in describing the behaviour of the entanglement entropy as the partition deviates from the event horizon. Inset shows the derivative of the entanglement entropy $S_{\mathcal{A}}(n_{h})$ with respect to $\ln(2n_{h})$ following a horizon partition, $\Delta n=0$, for increasing system sizes, indicating the regime where $N_{F}\approx 1$. As system size increases, the domain for which the gradient plateaus around 1 increases, suggesting that $N_{F}\rightarrow 1$ in the thermodynamic limit, $N\rightarrow\infty$.
@@ V.1 The breakdown of thermality
To further investigate the breakdown of thermality for non-horizon partitions, we examine the mode occupation expectation value, $\langle 0_{M}|c_{k}^{\dagger}c_{k}|0_{M}\rangle$, of the mean-field Hamiltonian that simulates the black hole space-time with respect to the ground state $|0_{M}\rangle$ of the asymptotically flat Minkowski Hamiltonian (Eq. (2) with couplings $v=0$ and $u=1$). Here, $c_{k}$ are the eigenmodes that diagonalise the mean-field Hamiltonian of Eq. (2) with the Gullstrand–Painlevé coupling profile $v(x)=\sqrt{x_{h}/x}$ and $u=1$, which correspond to a black hole space-time. The mode occupation can be expressed as [41]
@@
where $|E_{GP,k}\rangle$ and $|E_{M,k}\rangle$ denote the $k$th eigenstate of the single-particle Hamiltonians, $h_{GP}$ and $h_{M}$, associated with $H_{MF}$ for the Gullstrand-Painlevé and Minkowski coupling profiles, respectively. Their corresponding energy eigenvalues are denoted $E_{GP,k}$ and $E_{M,k}$. Evaluating the mode occupation $\langle 0_{M}|c_{k}^{\dagger}c_{k}|0_{M}\rangle$ is the lattice equivalent to determining the Minkowski vacuum expectation value perceived by an observer in a black hole space-time, which is the relevant quantity in the Unruh-Davies-Fulling and Hawking effects [41], and should not be interpreted as a dynamical quench.
To perform the partial trace of subsystem $\mathcal{A}$, we define the Gullstrand-Painlevé Hamiltonian only in the region corresponding to subsystem $\mathcal{B}$ and take it to be zero elsewhere. This effectively traces out the degrees of freedom of $\mathcal{A}$. As illustrated in Fig. 5, when the system is partitioned at the event horizon ($\Delta n=0$), the mode occupation expectation value of Eq. (23) follows a Fermi-Dirac distribution of the form
$f(E_{GP,k},T)=\frac{1}{e^{E_{GP,k}/T}+1},$ (24)
@@
Figure 5: Red data points show the mode occupation expectation value $\langle 0_{M}|c_{k}^{\dagger}c_{k}|0_{M}\rangle$ of the mean-field Hamiltonian that simulates the black hole space-time (Eq.(2) with the Gullstrand-Painlevé couplings $v(x)=\sqrt{x_{h}/x}$ and $u=1$, defined only on $\mathcal{B}$) with respect to the ground state $|0_{M}\rangle$ of the Minkowski Hamiltonian (Eq. (2) with couplings $u=1$ and $v=0$, defined over the entire lattice) for a horizon partition ($\Delta n=0$) and interior partition ($\Delta n=50$), respectively. The abscissa shows the energy eigenvalues $E_{GP,k}$ of the single-particle Hamiltonian $h_{GP}$ associated with $H_{MF}$ for the Gullstrand-Painlevé couplings; for finite NN the spectrum is bounded due to the lattice bandwidth. Blue curve shows data interpolated using the Fermi-Dirac distribution of Eq. (24) with $T=0.3271$ and $T=1.9237$, respectively. System sizes of $N=2000$ lattice sites, with the horizon located at $n_{h}=700$, were considered.
Figure 6: (a) Absolute value of interpolation errors $|e(E_{GP,k})|$ between the numerical and interpolated data for the mode occupation number $\langle 0_{M}|c_{k}^{\dagger}c_{k}|0_{M}\rangle$, interpolated using a Fermi-Dirac distribution, for systems of increasing $\Delta n$ with $N=2000$ lattice sites. As $\Delta n$ increases, and the system is partitioned further from $n_{h}$, increasing absolute interpolation errors indicate deviations from a Fermi-Dirac distribution in the mode occupation expectation value. (b) Integrated absolute interpolation errors $I_{e}(E_{GP,k})$ for increasing $\Delta n$ and varied system sizes NN, illustrating that the deviations from a Fermi-Dirac distribution as $\Delta n$ is increased are not minimised at larger system sizes. This is in contrast to partitioning the system at $n_{h}$, where the integrated interpolation errors decrease with increasing system size. For both figures, systems with $n_{h}=N/4$ were considered.
@@ V.2 Thermality and the Hawking temperature
We employ this method to probe the thermal properties of our chiral spin-chain simulator. However, the mean-field Hamiltonian of Eq. (2) does not feature an overall linear coupling and, therefore, does not approximate the entanglement Hamiltonian required to extract a thermal spectrum. To address this, the mean-field Hamiltonian must be modified such that the black hole geometry is encoded with only a linear nearest-neighbour coupling $u(x)$. In the continuum limit, such a modification corresponds to performing a coordinate transformation from Gullstrand–Painlevé to Schwarzschild coordinates, whilst preserving the space-time geometry (see Appendix D); thus, by mapping the black hole geometry to a mean-field Hamiltonian that approximates the entanglement Hamiltonian, the mode occupations of the partitioned ground state yields a Fermi-Dirac distribution from which $T_{H}$ can be reliably extracted [41, 20]. This procedure fails, however, when the chirality term $v(x)$ is non-zero, as in this case the mean-field Hamiltonian no longer approximates the corresponding entanglement Hamiltonian.
@@
which matches the Schwarzschild metric in $(1+1)$ dimensions. In particular, for the choice $u(x)=1-x_{h}/x$, this line element describes the exterior region of a Schwarzschild black hole with an event horizon located at $x_{h}$. This way, taking $v(x)$ to be zero and making $u(x)$ position dependent corresponds to a coordinate transformation from Gullstrand–Painlevé coordinates $(t,x)$ to Schwarzschild coordinates $(\tilde{t},x)$. Whilst the Hamiltonian’s structure changes under this transformation, scalar quantities, such as the space-time curvature, remain invariant (see Appendix D).
@@
According to the Fulling–Davies–Unruh effect [23, 19, 60], such an observer perceives the Minkowski vacuum as a thermal state with temperature $T_{H}=\kappa/2\pi$ [69] (see Appendices B). It follows that the partitioned ground state of the mean-field Hamiltonian of Eq. (2), with the linear coupling $u(x)$ of Eq. (26) and $v=0$, will be thermal with an effective temperature equal to the Hawking temperature $T_{H}$, as it approximates both the entanglement Hamiltonian and the Hamiltonian of a Dirac spinor in Rindler space-time [41, 20].
As in the Gullstrand–Painlevé case, the thermality of the Hamiltonian corresponding to a Dirac spinor in Rindler space-time can be confirmed by analysing the mode occupation of the partitioned system’s ground state. Specifically, we compute the expectation value of the mode occupation operators with respect to the ground state $|0_{M}\rangle$ of the Minkowski Hamiltonian. As shown in Fig. 7, the resulting mode occupations for the Rindler Hamiltonian follow a Fermi–Dirac distribution with temperature $T_{H}=\kappa/2\pi$, thereby confirming the thermal nature of the reduced state and validating the extraction of the Hawking temperature via this approach. Once again, this thermality is only present when the lattice system is partitioned at the horizon.
Figure 7: Data points show the mode occupation expectation values $\langle 0_{M}|c_{k}^{\dagger}c_{k}|0_{M}\rangle$ of the mean-field Hamiltonian that simulates Rindler space-time (Eq. (2) with couplings $u(x)=\alpha(x/x_{h}-1)$ and $v=0$, defined only on $\mathcal{B}$), computed with respect to the Minkowski ground state $|0_{M}\rangle$ (Eq. (1) with $u=1$ and $v=0$, defined over the entire lattice). The abscissa shows the energy eigenvalues $E_{R,k}$ of the single-particle Hamiltonian $h_{R}$ corresponding to $H_{MF}$ with the Rindler couplings. The blue curves correspond to a Fermi–Dirac distribution with temperature $T_{H}=\kappa/2\pi$, where $\kappa=u^{\prime}(x_{h})/2=\alpha/2x_{h}$ is the surface gravity associated with the linear coupling profile. System sizes $N=2000$, $2400$, $2800$, $3000$, and $3600$ were considered, with the horizon located at $n_{h}=N/2$ and coupling slope $\alpha=2$. Inset depicts the absolute interpolation error $|e(E_{R,k})|$ between numerical and interpolated data of the mode occupation for $\Delta n=0,20,40,60,80$ (colour scheme as in Fig. 6a) with $N=2000$ and $\alpha=1/2$.